MID-INFRARED DIODE-LASER DIAGNOSTICS OF THE GAS FLOW IN HIGH-MACH-NUMBER WIND TUNNELS

\underline{A.R. HIGHT WALKER}, G.T. FRASER, AND D.F. PLUSQUELLIC, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

A rapid-scan, lead-salt, diode-laser spectrometer has been constructed to investigate infrared active gases in high Mach number ( 8 - 14) wind tunnels for measurement of temporal temperature, pressure, and velocity profiles. The spectrometer is capable of capturing 4000 sweeps of a 0.2 cm^{−1} spectral window at a sampling rate of 10,000 sweeps/s with a minimum detectable fractional absorption of ∼ 1% per sweep. For room-temperature Doppler-limited absorption by CO, this corresponds to a detection sensitivity of ∼ 5 x 10^{−4} Torr-cm in a 100 μs. In addition to a spectral channel, the 12 bit data acquisition sampling system alternately captures an etalon trace to determine scan linearity and dispersion and a standard gas trace for correcting for frequency drifts. The choice of infrared chromophore to monitor is guided by real-time collection of a sample of the flow stream followed by laboratory analysis using Fourier-transform infrared spectroscopy. Testing of the spectrometer is presently being undertaken on a Mach 8 wind tunnel.